# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v1

@@ 3.9 Stochastic Gradient Descent - the Unstable Case
Proof of Theorem B (ii).
Suppose $x^{*}\in\mathcal{M}$ is regular with $\lambda(x^{*})>0$. In the following, we will show that there exists a $\delta>0$ such that the neighborhood $0\in V_{\delta}\subseteq\mathbb{R}^{D}$ from Lemma 3.2/Corollary 3.5 satisfies
$\mathbb{P}\left(\exists n\in\mathbb{N},\text{ s.t. }\psi_{\omega}^{(n)}(v,w)\notin V_{\delta}\right)=1,~{}\forall(v,w)\in V_{\delta}\setminus\left(\mathbb{R}^{D-N}\times\{0\}\right).$ (3.50)
@@
for all $(v,w)\in V_{\delta}\setminus(\mathbb{R}^{D-N}\times\{0\})$. ∎
@@ Acknowledgements
The authors would like to thank Tim van Erven for his insightful feedback. Furthermore, the authors thank the DFG SPP 2298 for supporting their research. Both authors have been additionally supported by Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689), in the case of D.C. via the Berlin Mathematical School and in the case of M.E. via projects AA1-8 and AA1-18. Furthermore, M.E. thanks the DFG CRC 1114 and the Einstein Foundation for support.
@@ Appendix A Comparison to second moment linear stability
A different notion of linear stability for SGD was introduced in [WME18]. Adapted to the notation introduced in section 3.5, their condition can be expressed as follows.
Definition A.1 ([WME18, Definition 2]).
@@
Second moment stability is closely related to so-called moment Lyapunov exponents (see e.g. [Arn84]), which have been linked to central limit theorems as well as large deviation theory for the convergence of finite-time Lyapunov exponents [BL85, AKO86, AK87].
Definition A.2.
@@
Proof.
Note that by (3.22), a global minimum $x^{*}$ is second moment linearly stable if and only if there exists a $C>0$ such that
@@
Proof.
As a consequence of Oseledet’s theorem [Ose68], we have
@@
The proposition follows from Jensen’s inequality. ∎
To summarize, if $x^{*}\in\mathcal{M}$ is second moment stable in the sense of [WME18], it satisfies $\Lambda_{2}(x^{*})\leq 0$ and thus $\lambda(x^{*})\leq 0$. Most of these points should be regular in the sense of Definition 2.1 and even satisfy $\lambda(x^{*})<0$. By Theorem B, these global minima lie in the support of $X_{\lim}^{\operatorname{SGD}}$. On the other hand the one dimensional example demonstrated above suggests that it is possible to have global minima $x^{*}\in\mathcal{M}$ with $\lambda(x^{*})<0$ and $\Lambda_{2}(x^{*})>0$. Given that they are regular, these global minima will be in the support of $X_{\lim}^{\operatorname{SGD}}$, but will not be second moment linearly stable in the sense of [WME18]. Notably [WME18] contains a sufficient condition for second order linear stability (cf. [WME18, Theorem 1]). By the argument above, their condition is also a sufficient condition for $\lambda(x^{*})\leq 0$.
In [MY21], a more general notion of “$k$-th order linear stability” is introduced. If $p=k$ is an even natural number this is equivalent to the following generalization of second moment linear stability (cf. [MY21, Remark 1]).
Definition A.5.

### v2

@@ 3.9 Stochastic Gradient Descent - the Unstable Case
Proof [of Theorem B (ii)] Suppose $x^{*}\in\mathcal{M}$ is regular with $\lambda(x^{*})>0$. In the following, we will show that there exists a $\delta>0$ such that the neighborhood $0\in V_{\delta}\subseteq\mathbb{R}^{D}$ from Lemma 3.2/Corollary 3.5 satisfies
$\mathbb{P}\left(\exists n\,\in\mathbb{N},\text{ s.t. }\psi_{\omega}^{(n)}(v,w)\notin V_{\delta}\right)=1,~{}\forall,(v,w)\in V_{\delta}\setminus\left(\mathbb{R}^{D-N}\times\{0\}\right).$ (3.50)
@@
for all $(v,w)\in V_{\delta}\setminus(\mathbb{R}^{D-N}\times\{0\})$.
@@ Acknowledgements
The authors would like to thank Tim van Erven for his insightful feedback. Furthermore, the authors thank the DFG SPP 2298 for supporting their research. Both authors have been additionally supported by Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689), in the case of D.C. via the Berlin Mathematical School and in the case of M.E. via projects AA1-8, AA1-18 and EF45-5. Furthermore, M.E. thanks the DFG CRC 1114 and the Einstein Foundation for support.
@@ Appendix A Comparison to second moment linear stability
A different notion of linear stability for SGD was introduced by Wu et al. (2018). Adapted to the notation introduced in section 3.5, their condition can be expressed as follows.
Definition A.1 (Definition 2 in Wu et al. (2018))
@@
Second moment stability is closely related to so-called moment Lyapunov exponents (see e.g. Arnold (1984)), which have been linked to central limit theorems as well as large deviation theory for the convergence of finite-time Lyapunov exponents (see e.g. Bougerol and Lacroix (1985); Arnold et al. (1986); Arnold and Kliemann (1987)).
Definition A.2
@@
Proof  Note that by (3.22), a global minimum $x^{*}$ is second moment linearly stable if and only if there exists a $C>0$ such that
@@
Proof  As a consequence of Oseledec theorem (cf.  Theorem 2.1), we have
@@
The proposition follows from Jensen’s inequality.
To summarize, if $x^{*}\in\mathcal{M}$ is second moment stable in the sense of Wu et al. (2018), it satisfies $\Lambda_{2}(x^{*})\leq 0$ and thus $\lambda(x^{*})\leq 0$. Most of these points should be regular in the sense of Definition 2.2 and even satisfy $\lambda(x^{*})<0$. By Theorem B, these global minima lie in the support of $X_{\lim}^{\operatorname{SGD}}$. On the other hand the one dimensional example demonstrated above suggests that it is possible to have global minima $x^{*}\in\mathcal{M}$ with $\lambda(x^{*})<0$ and $\Lambda_{2}(x^{*})>0$. Given that they are regular, these global minima will be in the support of $X_{\lim}^{\operatorname{SGD}}$, but will not be second moment linearly stable in the sense of Wu et al. (2018). Notably the paper by Wu et al. (2018) contains a sufficient condition for second order linear stability (cf.  Therorem 1 in Wu et al. (2018)). By the argument above, their condition is also a sufficient condition for $\lambda(x^{*})\leq 0$.
In Ma and Ying (2021), a more general notion of “$k$-th order linear stability” is introduced. If $p=k$ is an even natural number this is equivalent to the following generalization of second moment linear stability (cf. Remark 1 in Ma and Ying (2021)).
Definition A.5